OpenStudy (anonymous):

Solve 3x^2 + 4x = 2.

OpenStudy (turingtest):

3x^2+4x-2=0 quadratic formula from here. You know it, right?

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

apply quadratic formula first of all right it in standard form as 3x\[^{2}\]+4x-2=0 then write finally u get \[x=(-2\pm \sqrt{10})\div\]

OpenStudy (anonymous):

where is my full answer:*(

OpenStudy (anonymous):

It cut you off :(

OpenStudy (anonymous):

wait i will right again

OpenStudy (anonymous):

is it divided my 3?

OpenStudy (turingtest):

\[x=(-b \pm \sqrt{b^2-4ac})/2a\]

OpenStudy (anonymous):

*by

OpenStudy (anonymous):

yes diviided by 3

OpenStudy (turingtest):

where the quadratic is in the form ax^2+bx+c=0

OpenStudy (anonymous):

3x^2+4x-2=0 \[x=(-b \pm \sqrt{b^2 -4ac})/2a\] \[x=(-4 \pm \sqrt{4^2 -4(3)(-2)})/2(3)\] \[x=(-4 \pm 2\sqrt{10})/6\]\[x=(-4+2\sqrt{10})/6\]or\[x=(-4-2\sqrt{10})/6\]

OpenStudy (anonymous):

Oh okay well thank you :D

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

simplify it by calculator